The Module5QuizData_HudsonHighSchool.xlsx spreadsheet contains the income of randomly selected families whose students attend Hudson and Cedar Falls High Schools. You want to test Ho: Hudson mean income = Cedar Falls mean income against Ha: Hudson mean income ≠ Cedar Falls mean income. For α = 0.05, what do you conclude?

a) We reject Ho, because the p-value is 0.02.

b) We reject Ho, because the p-value is 0.04.

c) We fail to reject Ho, because the p-value is 0.12.

d) We fail to reject Ho, because the p-value is 0.24.
Answer #1

First we need to calculate mean and standard deviation for Hudson and Cedar Falls High Schools.

Hudson School

Mean = 54,865.087
Standard Deviation = 9093.7446

Cedar Falls School
Mean = 52,080.250
Standard Deviation = 7631.0978

Test Statistics

Degrees of Freedom = n1 + n2 - 2

Degrees of Freedom df = 23 + 28 - 2 = 49

P-value

P-value corresponding to t= 1.189 and df = 49 is 0.24 (Obtained using online calculator. Screenshot attached)

Decision about the null hypothesis

Since p-value = 0.24 is greater than α = 0.05, we fail to reject null hypothesis Ho.

Option D is correct